# Large mechanical squeezing through synthetic magnetism in optomechanics

### v1

@@ Abstract
We propose a scheme to generate large amount of mechanical squeezing, far beyond the $\rm{3dB}$ limit, which is based on synthetic magnetism in optomechanical system that hosts a Backward Stimulated Brillouin Scattering (BSBS) process. Our benckmark system consists of an acoustic mode coupled to two optical modes through the BSBS process, and a Duffing mechanical oscillator that couples to the same optical modes through the standard optomechanical radiation pressure. The synthetic magnetism comes from the modulation of the mechanical coupling between the acoustic and the mechanical modes. When there is no synthetic magnetism, a given amount of mechanical squeezing is generated in the system. This squeezing is mainly dependent on the BSBS process, and it is fragile against thermal noise. By switching on the synthetic magnetism, the degree of the generated squeezing is greatly enhanced and goes far beyond the limit of the $\rm{3dB}$. This large magnetism induced squeezing persists even when there is no BSBS process in the system. Moreover, this generated squeezing is robust enough against thermal noise in comparison to the one induced when the synthetic magnetism is off. Furthermore, both the mechanical variance squeezing and effective phonon number exhibit series of peaks and dips depending on the phase modulation of the mechanical coupling. This oscillatory feature is reminscent of a sudden death and revival of squeezing phenomenon, which can be used to maintain a desired magnitude of squeezing by tuning this phase. Our proposal provides a path toward a flexible scheme that generates large amount of squeezing, far beyond the $\rm{3dB}$ limit. Such a generated squeezed states can be use for quantum applications including quantum information processing, quantum sensing and metrology, and quantum computing.
Optomechanics, synthetic magnetism, squeezing
@@ I Introduction
Nonclassical states such as squeezed and entangled states are crucial ingredients required to improve a range of quantum applications including quantum information processing Wendin (2017); Slussarenko and Pryde (2019); Meher and Sivakumar (2022), quantum sensing/metrology Degen et al. (2017); Polino et al. (2020); Stray et al. (2022); Barbieri (2022), quantum computing Zidan et al. (2021), and quantum supremacy Arute et al. (2019); Zhong et al. (2020, 2021). For instance, entangled states have been widely generated in optomechanical systems Aspelmeyer et al. (2014); Foulla et al. (2017) by exploring diverse techniques Riedinger et al. (2018); Tchodimou et al. (2017); Kotler et al. (2021); Chakraborty and Sarma (2019); Massembele et al. (2024), and they were recently proposed as resources to enhance sensing Xia et al. (2023); Brady et al. (2023). Beside that, large intracavity squeezed field have been generated Vahlbruch et al. (2016); Qin et al. (2022), which can be used for mechanical cooling Clark et al. (2017); Djorwé et al. (2012), gravitational wave detection Aasi et al. (2013); et al. (2016), and to further generate nonclassical states Qin et al. (2021). Similarly, strong mechanical squeezed states Banerjee et al. (2023) were equally generated, which can be useful to enhance mass-sensing Djorwe et al. (2019); Tchounda et al. (2023); Djorwé et al. (2023a), phonon information processing Stannigel et al. (2012); Madiot et al. (2023), and for phononic state transport/transfer Ren et al. (2022). Despite these interesting applications related to squeezed states, their generation is often limited by quantum noise that resists any measurement below the Zero Point Fluctuation (ZPF), i.e., beyond the $\rm{3dB}$ limit. The existing systems in which mechanical squeezing beyond $\rm{3dB}$ has been achieved have explored reservoir engineering technique or two-tone driving Woolley and Clerk (2014); Wollman et al. (2015); Lei et al. (2016).
@@
Recently, technique based on synthetic magnetism (engineered via a modulation of the photon/phonon hopping rate) have been used in optomechanical structures for specific purposes. For instance, an artificial magnetic field for photons was engineered to achieve photon transport in Schmidt et al. (2015); Fang et al. (2012). Similarly, synthetic magnetic field for phonon/acoustic has been created in Brendel et al. (2017); Mathew et al. (2020); Wang et al. (2020) for phononic transport at nanoscale. More recently, synthetic magnetism has been engineered in optomechanical systems to enhance entangled states generation Lai et al. (2022). Owing to these interesting physics fostered by synthetic magnetism in optomechanics, here we use it to enhance mechanical squeezing in optomechanical system involving Backward Stimulated Brillouin Scattering (BSBS) process which has been recently proposed Chen et al. (2023).
@@
The underligned system consists of an acoustic (mechanical) mode coupled to two optical modes through the BSBS process (radiation pressure coupling). The merit of our proposal is to connect the acoustic and the mechanic modes through a mechanical coupling having a strengh $J_{m}$ that is modulated via a phase $\theta$. Such a phase modulation of the coupling induces a synthetic magnetism in our benckmark system Lai et al. (2022); Djorwé et al. (2023b). When the synthetic magnetism is off, a given amount of mechanical squeezing is generated in the system and this squeezing is mainly dependent on the BSBS process, and it is fragile against thermal noise Chen et al. (2023). By switching on the synthetic magnetism, we found that i) the degree of the generated squeezing is greatly enhanced and goes far beyond the limit of the $\rm{3dB}$, ii) this squeezing persists even when there is no BSBS process in the system, and iii) this generated squeezing is robust enough against thermal noise in comparison to the case when the synthetic magnetism is off. Furthermore, both the mechanical squeezing and effective phonon number exhibit series of peaks and dips depending on the phase modulation of the mechanical coupling. This oscillatory feature induces a sudden death and revival of squeezing, which can be used to maintain a desired magnitude of squeezing by tuning this phase. Our proposal pave a way toward a flexible scheme that can be used to generate arbitrary amount of squeezing. Such a generated squeezed states can be used for quantum applications including quantum information processing, quantum sensing and metrology, and quantum computing.
@@ II Model and dynamical equations
Our system consists of an acoustic mode ($b_{a}$) that couples to two optical modes through the BSBS process, and a nonlinear mechanical oscillator ($b_{m}$) that couples also to the two optical modes. In order to enhance the squeezing of the mechanical oscillator, we assume that the acoustic and mechanical modes are mechanically coupled through a coupling strengh $J_{m}$, which is modulated through a phase $\theta$ that induces a synthetic jauge into the dynamics. The Hamiltonian of such a system is given by ($\hbar=1$):
@@
$H_{\rm{OM}}$ ::=:= $g_{m}\sum_{j=1,2}a_{j}^{\dagger}a_{j}(b_{m}+b_{m}^{\dagger})+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4},$ (3)
@@
In the above Hamiltonian, the first term $H_{\rm{0}}$ is the free Hamiltonian of the optical modes $a_{j}$, the acoustic and for the mechanical mode. The first term in $H_{\rm{OM}}$ capures the optomechanical interaction between the optical and the mechanical modes, while the second one acconts for nonlinear effect on the mechanical resonator through the Duffing coefficient $\eta$. The third term $H_{\rm{BSBS}}$ stands for the triply resonant phonon-photon interaction striggered via the $\rm{BSBS}$ process, and $H_{\rm{int}}$ depicts the mechanical coupling between the acoustic ($b_{a}$) and the mechanical ($b_{m}$) modes. The single-photon optomechanical (Brillouin) coupling is captured by $g_{m}$ ($g_{a}$), which comes from the radiation pressure (electrostrictive) force. The drivings fields are captured by the last term $H_{\rm{drive}}$, where $E_{j}$ and $\omega_{p_{j}}$ are the amplitude and frequency of the $j^{th}$ field. In what follow, we will assume the same frequency field, $\omega_{p_{1,2}}\equiv\omega_{p}$. The optical cavity frequency are $\omega_{c_{j}}$ and the mechanical (acoustic) frequency is $\omega_{m}$ ($\omega_{a}$). In the frame rotating at $\frac{\omega_{p}}{2}a_{1}^{\dagger}a_{1}+\omega_{p}a_{2}^{\dagger}a_{2}+\frac{\omega_{p}}{2}b_{a}^{\dagger}b_{a}$, the Hamiltonian in Equation 1 becomes
$H$ $=-\Delta a_{1}^{\dagger}a_{1}+\Delta_{a}b_{a}^{\dagger}b_{a}+\omega_{m}b_{m}^{\dagger}b_{m}-g_{m}a_{1}^{\dagger}a_{1}(b_{m}+b_{m}^{\dagger})$ (7)
$+{J_{m}}({e^{i\theta}}{b_{a}^{\dagger}}{b_{m}}+{e^{-i\theta}}{b_{a}}{b_{m}^{\dagger}})+iE_{l_{1}}(a_{1}^{\dagger}-a_{1})$
$+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4}-{G_{a}}(a_{1}^{\dagger}b_{a}+a_{1}b_{a}^{\dagger}),$
where we have defined $\Delta=\omega_{l_{1}}-\omega_{c_{1}}$, $\Delta_{a}=\omega_{a}-\omega_{l_{1}}$ and ${G_{a}}=g_{a}\alpha_{2}$. Here $\alpha_{2}$ is the steady-state of the control optical mode $a_{2}$, which has been treated classically as it is assumed strong compared to the weak strengh of the Brillouin acoustic mode $b_{a}$.
Figure 1: Sketch of our linearized three mode optomechanical system. The mode $a_{1}$ is coupled to the acoustic (mechanical) mode $b_{a}$ ($b_{m}$) the Brillouin (optomechanical) coupling $G_{a}$ ($G_{a}$), which are induced from electrostrictive (radiation pressure) force. The phonon-phonon hopping rate $J_{m}$ is modulated by the phase $\theta$.
By following the standard linearization process of optomechanical process, the Hamiltonian in Equation 7 can be linearized, and that reduces our proposal to a three modes optomechanical system sketched as in Figure 1. To investigate on the squeezing of the target mechanical resonator $b_{m}$, we introduce the following Bogoliubov transformation,
@@
$-(G^{\prime}_{m}\delta a_{1}^{\dagger}+G_{m}^{{}^{\prime}*}\delta a_{1})(\delta b_{m}^{s\dagger}+\delta b_{m}^{s})+{J^{\prime}_{m}}({e^{i\theta}}{\delta b_{a}^{\dagger}}\delta b_{m}^{s}$
@@
with the effective parameters $\omega^{\prime}_{m}=\sqrt{\omega_{m}{(\omega_{m}+2\Lambda)}}$, $G^{\prime}_{m}=G_{m}e^{-r}$ and $J^{\prime}_{m}=J_{m}\cosh(r)$, and the effective mechanical coupling is $G_{m}=g_{m}\alpha_{1}$, where $\Lambda=24\eta\Re(\beta_{m})^{2}$ by keeping in mind that $\beta_{m}$ ($\alpha_{1}$) is the steady-state of the mechanical (optical) mode. By taking into account the optical ($\kappa$), mechanical ($\gamma_{m}$) and acoustic ($\gamma_{a}$) dissipations, one can write the dynamics of the fluctuations operators as,
$\begin{cases}\delta\dot{a}_{1}&=\left(i\tilde{\Delta}-\frac{\kappa}{2}\right)\delta a_{1}+iG^{\prime}_{c}({\delta}b_{m}^{s\dagger}+{\delta}b_{m}^{s})+i{G_{b}}\delta b_{a}\\ &+\sqrt{\kappa}a_{1}^{in}\\ \delta\dot{b}_{a}&=-({\frac{\gamma_{a}}{2}}+i{\Delta_{a}})\delta{b}_{a}-i{J^{\prime}_{m}}e^{i\theta}\delta{b}_{m}^{s}+i{G_{a}}\delta a_{1}\\ &+\sqrt{\gamma_{a}}b_{a}^{in}\\ \delta\dot{b}_{m}^{s}&=-({\frac{\gamma_{m}}{2}}+i{\omega^{\prime}_{m}})\delta{b}_{m}^{s}-i{J^{\prime}_{m}}e^{-i\theta}\delta{b}_{a}+i(G_{m}^{{}^{\prime*}}\delta a_{1}\\ &+G_{m}^{\prime}\delta a_{1}^{\dagger})+\sqrt{\gamma_{m}}b_{m}^{s^{in}},\end{cases}$ (10)
@@
$\langle a_{1}^{in}(t)a_{1}^{in\dagger}(t^{\prime})\rangle=\delta(t-t^{\prime}),\hskip 14.22636pt\langle a_{1}^{in\dagger}(t)a_{1}^{in}(t^{\prime})\rangle=0,$
@@
$\langle b_{m}^{in}(t)b_{m}^{in}(t^{\prime})\rangle=(n_{th}+\frac{1}{2})\sinh(2r)\delta(t-t^{\prime}),$
where $n_{th}$ is the equilibrium phonon occupation of the mechanical resonator defined as $n_{th}=[\rm exp(\frac{\hbar\omega_{m}}{k_{b}T})-1]^{-1}$, where $\rm k_{b}$ is the Boltzmann constant. Owing to the high-frequency Brillouin mode $b_{a}$ compared to the mechanical one ($\omega_{m}\ll\omega_{a}$) the thermal acoustic occupancy has been neglected. In order to figure out the effect of the mechanical coupling $J_{m}$ and its phase modulation $\theta$ on the squeezing degree of the targeted mechanical resonator, we start by defined the following amplitude (position) and phase (momentum) quadrature operators, $\delta X_{I,q_{a,m}}=\frac{\delta\mathcal{O}^{\dagger}+\delta\mathcal{O}}{\sqrt{2}}$, $\delta Y_{\phi,p_{a,m}}=i\frac{\delta\mathcal{O}^{\dagger}-\delta\mathcal{O}}{\sqrt{2}}$, with $\mathcal{O}\equiv a_{1},b_{a},b_{m}^{s}$. Similarly, the related noise quadratures read, $\delta X_{I,q_{a,m}}^{in}=\frac{\delta\mathcal{O}^{\dagger in}+\delta\mathcal{O}^{in}}{\sqrt{2}}$, $\delta Y_{\phi,p_{a,m}}^{in}=i\frac{\delta\mathcal{O}^{\dagger in}-\delta\mathcal{O}^{in}}{\sqrt{2}}$. Therefore, one derives the set of equations describing the quadrature dynamics of our system as,
$\dot{u}={\rm M}u+z^{in},$ (12)
where ${u}=(\delta I,\delta\phi,\delta q_{a},\delta p_{a},\delta q_{m},\delta p_{m})^{T}$, $z^{in}=(\sqrt{\kappa}I^{in},\sqrt{\kappa}\phi^{in},\sqrt{\gamma_{a}}q_{a}^{in},\sqrt{\gamma_{a}}p_{a}^{in},\sqrt{\gamma_{m}}q_{m}^{in},\sqrt{\gamma_{m}}p_{m}^{in},)^{T}$ and the matrix $\rm M$ is given by,
${\rm M}=\begin{pmatrix}-\frac{\kappa_{1}}{2}&-\tilde{\Delta}&0&-G_{a}&0&0\\ \tilde{\Delta}&-\frac{\kappa_{1}}{2}&G_{a}&0&2G_{m}^{{}^{\prime}}&0\\ 0&-G_{a}&-\frac{\gamma_{a}}{2}&\Delta_{a}^{{}^{\prime}}&J^{\prime}_{m}\sin{\theta}&J^{\prime}_{m}\cos{\theta}\\ G_{a}&0&-\Delta_{a}^{{}^{\prime}}&-\frac{\gamma_{a}}{2}&-J^{\prime}_{m}\cos{\theta}&J^{\prime}_{m}\sin{\theta}\\ 0&0&-J^{\prime}_{m}\sin{\theta}&J^{\prime}_{m}\cos{\theta}&-\frac{\gamma_{m}}{2}&\omega_{m}^{{}^{\prime}}\\ 2G_{m}^{{}^{\prime}}&0&-J^{\prime}_{m}\cos{\theta}&-J^{\prime}_{m}\sin{\theta}&-\omega_{m}^{{}^{\prime}}&-\frac{\gamma_{m}}{2}\end{pmatrix},$ (13)
where the effective couplings $G_{m}^{{}^{\prime}}$ and $G_{a}$ have been assumed real for simplicity.

### v2

# Synthetic magnetism enhanced mechanical squeezing in Brillouin optomechanical system

@@ Abstract
We propose a scheme to generate large amount of mechanical squeezing, far beyond the $\rm{3dB}$ limit, which is based on synthetic magnetism in optomechanical system that hosts a Backward Stimulated Brillouin Scattering (BSBS) process. Our benchmark system consists of an acoustic mode coupled to two optical modes through the BSBS process, and a Duffing mechanical oscillator that couples to the same optical modes through the standard optomechanical radiation pressure. The synthetic magnetism comes from the modulation of the mechanical coupling between the acoustic and the mechanical mode. When there is no synthetic magnetism, a given amount of mechanical squeezing is generated in the system. This squeezing is mainly dependent on the BSBS process, and it is fragile against thermal noise. By switching on the synthetic magnetism, the degree of the generated squeezing is greatly enhanced and goes far beyond the limit of the $\rm{3dB}$. This large magnetism induced squeezing persists even when there is no BSBS process in the system. Moreover, this generated squeezing is robust enough against thermal noise in comparison to the one induced when the synthetic magnetism is off. Furthermore, both the mechanical variance squeezing and effective phonon number exhibit series of peaks and dips depending on the phase modulation of the mechanical coupling. This oscillatory feature is reminiscent of a sudden death and revival of squeezing phenomenon, which can be used to maintain a desired magnitude of squeezing by tuning this phase. Our proposal provides a path toward a flexible scheme that generates large amount of squeezing, far beyond the $\rm{3dB}$ limit. Such a generated squeezed states can be used for quantum applications including quantum information processing, quantum sensing and metrology, and quantum computing.
Optomechanics, synthetic magnetism, squeezing, Brillouin Scattering
@@ I Introduction
Nonclassical states such as squeezed and entangled states are crucial ingredients required to improve a range of quantum applications including quantum information processing Wendin (2017); Slussarenko and Pryde (2019); Meher and Sivakumar (2022), quantum sensing and metrology Degen et al. (2017); Polino et al. (2020); Stray et al. (2022); Barbieri (2022), quantum computing Zidan et al. (2021), and quantum supremacy Arute et al. (2019); Zhong et al. (2020, 2021). For instance, entangled states have been widely generated in optomechanical systems Aspelmeyer et al. (2014); Foulla et al. (2017) by exploring diverse techniques Riedinger et al. (2018); Tchodimou et al. (2017); Kotler et al. (2021); Chakraborty and Sarma (2019); Massembele et al. (2024), and they were recently proposed as resources to enhance sensing Xia et al. (2023); Brady et al. (2023). Beside that, large intracavity squeezed field have been generated Vahlbruch et al. (2016); Qin et al. (2022), which can be used for mechanical cooling Clark et al. (2017); Djorwé et al. (2012), gravitational wave detection Aasi et al. (2013); et al. (2016), and to further generate nonclassical states Qin et al. (2021). Similarly, strong mechanical squeezed states Banerjee et al. (2023) were equally generated, which can be useful to enhance mass-sensing Djorwe et al. (2019); Tchounda et al. (2023); Djorwé et al. (2023a), information processing Stannigel et al. (2012); Madiot et al. (2023), and state transfer Ren et al. (2022) involving phonons. Despite these interesting applications related to squeezed states, their generation is often limited by quantum noise that resists any measurement below the Zero Point Fluctuation (ZPF), i.e., beyond the $\rm{3dB}$ limit. The existing systems in which mechanical squeezing beyond $\rm{3dB}$ has been achieved have explored reservoir engineering technique or two-tone driving Woolley and Clerk (2014); Wollman et al. (2015); Lei et al. (2016).
@@
Recently, technique based on synthetic magnetism (engineered via a modulation of the photon/phonon hopping rate) have been used in optomechanical structures for specific purposes. For instance, an artificial magnetic field for photons was engineered to achieve photon transport as reported in Schmidt et al. (2015); Fang et al. (2012). Similarly, synthetic magnetic field for phonon/acoustic has been created Brendel et al. (2017); Mathew et al. (2020); Wang et al. (2020) for phononic transport at the nanoscale. More recently, synthetic magnetism has been engineered in optomechanical systems to enhance entangled states generation Lai et al. (2022). Owing to these interesting physics fostered by synthetic magnetism in optomechanics, here we use it to enhance mechanical squeezing in optomechanical system involving Backward Stimulated Brillouin Scattering (BSBS) process which has been proposed in Shen et al. (2021).
@@
The underlying system, we are considering, consists of an acoustic (mechanical) mode coupled to two optical modes through the BSBS process (radiation pressure coupling). We propose a scheme to connect the acoustic and the mechanic modes through a mechanical coupling having a strength $J_{m}$ that is modulated via a phase $\theta$. Such a phase modulation of the coupling induces a synthetic magnetism in our model system Lai et al. (2022); Djorwé et al. (2023b). When the synthetic magnetism is switched off, a given amount of mechanical squeezing is generated in the system. This squeezing is mainly dependent on the BSBS process, and it is fragile against thermal noise Shen et al. (2021). By switching on the synthetic magnetism, we found that i) the degree of the generated squeezing is greatly enhanced and goes far beyond the limit of the $\rm{3dB}$, ii) this squeezing persists even when there is no BSBS process in the system, and iii) this generated squeezing is robust enough against thermal noise in comparison to the case when the synthetic magnetism is off. Furthermore, both the mechanical squeezing and effective phonon number exhibit series of peaks and dips depending on the phase modulation of the mechanical coupling. This oscillatory feature induces a sudden death and revival of squeezing, which can be used to maintain a desired magnitude of squeezing by tuning the phase. Our proposal paves a way towards a flexible scheme that can be used to generate arbitrary amount of squeezing. Such a generated squeezed states can be used for quantum applications including quantum information processing, quantum sensing/metrology, and quantum computing.
@@ II Model and dynamical equations
We consider a system consisting of an acoustic mode ($b_{a}$) that couples to two optical modes through the BSBS process, and a nonlinear mechanical oscillator ($b_{m}$) that couples also to the two optical modes. In order to enhance the squeezing of the mechanical oscillator, the acoustic and the mechanical modes are mechanically coupled, having coupling strength $J_{m}$. The coupling is modulated through a phase $\theta$ that induces a synthetic gauge into the dynamics. The Hamiltonian of such a system is given by ($\hbar=1$):
@@
$H_{\rm{OM}}$ ::=:= $\sum_{j=1,2}g_{a,m}a_{j}^{\dagger}a_{j}(b_{m}+b_{m}^{\dagger})+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4},$ (3)
@@
In the above Hamiltonian, the first term $H_{\rm{0}}$ is the free Hamiltonian corresponding to the optical ($a_{j}$), the acoustic ($b_{a}$) and the mechanical ($b_{m}$) mode. The first term in $H_{\rm{OM}}$ captures the optomechanical interaction between the optical and the mechanical modes, while the second one accounts for nonlinear effect on the mechanical resonator through the Duffing coefficient $\eta$. The third term $H_{\rm{BSBS}}$ stands for the triply resonant phonon-photon interaction triggered via the $\rm{BSBS}$ process. $H_{\rm{int}}$ describes the mechanical coupling between the acoustic and the mechanical modes. The single-photon optomechanical (Brillouin) coupling, $g_{m}$ ($g_{a}$), results from the radiation pressure (electrostrictive) force. The last term, $H_{\rm{drive}}$, denotes the drivings fields. Here $E_{j}$ and $\omega_{p_{j}}$ are the amplitude and the frequency of the $j^{th}$ field. In what follows, we will assume the same frequency field, $\omega_{p_{1,2}}\equiv\omega_{p}$. The optical cavity frequency are $\omega_{c_{j}}$ and the mechanical (acoustic) frequency is $\omega_{m}$ ($\omega_{a}$). In the frame rotating at $H_{r}=\omega_{p_{1}}a_{1}^{\dagger}a_{1}+\omega_{p_{2}}a_{2}^{\dagger}a_{2}+({\omega_{p_{1}}-\omega_{p_{2}}})b_{a}^{\dagger}b_{a}$, the Hamiltonian in Equation 1 becomes
$H^{\prime}$ $=-\Delta_{1}a_{1}^{\dagger}a_{1}+\Delta_{a}b_{a}^{\dagger}b_{a}+\omega_{m}b_{m}^{\dagger}b_{m}-g_{m}a_{1}^{\dagger}a_{1}(b_{m}+b_{m}^{\dagger})$ (7)
$+{J_{m}}({e^{i\theta}}{b_{a}^{\dagger}}{b_{m}}+{e^{-i\theta}}{b_{a}}{b_{m}^{\dagger}})+\frac{\eta}{2}(b_{m}+b_{m}^{\dagger})^{4}$
$+iE_{1}(a_{1}^{\dagger}-a_{1})-G_{a}(a_{1}^{\dagger}b_{a}+a_{1}b_{a}^{\dagger}),$
where we have defined $\Delta_{1}=\omega_{p_{1}}-\omega_{c_{1}}$, and $\Delta_{a}=\omega_{a}+\omega_{p_{2}}-\omega_{p_{1}}$. Here $\alpha_{2}$ is the steady-state of the control optical mode $a_{2}$, which has been treated classically as it is assumed to be strong compared to the weak strength of the Brillouin acoustic mode $b_{a}$ (see details in Appendix A).
Figure 1: Sketch of our linearized three mode optomechanical system. The mode $a_{1}$ is coupled to the acoustic ($b_{a}$) and the mechanical mode ($b_{m}$) with coupling strength $G_{a}$ and $G_{m}$ respectively. The mechanical oscillator has the Duffing coefficient $\eta$. The phonon-phonon hopping rate $J_{m}$ is modulated by the phase $\theta$.
By following the standard linearization procedure (Appendix A), the Hamiltonian in Equation 7 can be linearized, and that reduces our scheme to a three-mode optomechanical system as depicted in Figure 1. In order to investigate the squeezing of the mechanical resonator $b_{m}$, we introduce the following Bogoliubov transformation,
@@
$-(G^{\prime}_{m}\delta a_{1}^{\dagger}+G_{m}^{{}^{\prime}\ast}\delta a_{1})(\delta b_{m}^{s\dagger}+\delta b_{m}^{s})+{J^{\prime}_{m}}({e^{i\theta}}{\delta b_{a}^{\dagger}}\delta b_{m}^{s}$
@@
with the effective parameters $\omega^{\prime}_{m}=\sqrt{\omega_{m}{(\omega_{m}+2\Lambda)}}$, $G^{\prime}_{m}=G_{m}e^{-r}$ and $J^{\prime}_{m}=J_{m}\cosh(r)$. The effective mechanical coupling is $G_{m}=g_{m}\alpha_{1}$, where $\Lambda=24\eta\Re(\beta_{m})^{2}$. $\beta_{m}$ ($\alpha_{1}$) is the steady-state of the mechanical (optical) mode.
By taking into account the dissipation associated with the optical ($\kappa$), the mechanical ($\gamma_{m}$) and the acoustic ($\gamma_{a}$) modes, one can obtain the following equations describing the dynamics of the fluctuations operators,
$\begin{cases}\delta\dot{a}_{1}&=\left(i\tilde{\Delta}-\frac{\kappa}{2}\right)\delta a_{1}+iG^{\prime}_{m}({\delta}b_{m}^{s\dagger}+{\delta}b_{m}^{s})+i{G_{a}}\delta b_{a}\\ &+\sqrt{\kappa}a_{1}^{in}\\ \delta\dot{b}_{a}&=-({\frac{\gamma_{a}}{2}}+i{\Delta_{a}})\delta{b}_{a}-i{J^{\prime}_{m}}e^{i\theta}\delta{b}_{m}^{s}+i{G_{a}}\delta a_{1}\\ &+\sqrt{\gamma_{a}}b_{a}^{in}\\ \delta\dot{b}_{m}^{s}&=-({\frac{\gamma_{m}}{2}}+i{\omega^{\prime}_{m}})\delta{b}_{m}^{s}-i{J^{\prime}_{m}}e^{-i\theta}\delta{b}_{a}+i(G_{m}^{{}^{\prime}\ast}\delta a_{1}\\ &+G_{m}^{\prime}\delta a_{1}^{\dagger})+\sqrt{\gamma_{m}}b_{m}^{s^{in}},\end{cases}$ (10)
@@
$\langle a_{1}^{in}(t)a_{1}^{in\dagger}(t^{\prime})\rangle$ $=\delta(t-t^{\prime}),\hskip 14.22636pt\langle a_{1}^{in\dagger}(t)a_{1}^{in}(t^{\prime})\rangle=0,$ (11)
@@
$\langle b_{m}^{in}(t)b_{m}^{in}(t^{\prime})\rangle$ $=(n_{th}+\frac{1}{2})\sinh(2r)\delta(t-t^{\prime}).$
Here, $n_{th}$ represents the equilibrium phonon occupation number of the mechanical resonator, and is defined as, $n_{th}=[\rm exp(\frac{\hbar\omega_{m}}{k_{b}T})-1]^{-1}$, where $\rm k_{b}$ is the Boltzmann constant. Owing to the high-frequency acoustic mode $b_{a}$ compared to the mechanical one ($\omega_{m}\ll\omega_{a}$) the thermal acoustic phonon occupation has been neglected. In order to investigate the effect of the mechanical coupling $J_{m}$ and its phase modulation $\theta$ on squeezing of the targeted mechanical resonator, we first define the following amplitude (position) and phase (momentum) quadrature operators: $\delta X_{\mathcal{O}}=\frac{\delta\mathcal{O}^{\dagger}+\delta\mathcal{O}}{\sqrt{2}}$, $\delta Y_{\mathcal{O}}=i\frac{\delta\mathcal{O}^{\dagger}-\delta\mathcal{O}}{\sqrt{2}}$, with $\mathcal{O}\equiv a_{1},b_{a},b_{m}^{s}$. Similarly, the related noise quadratures read as: $\delta X_{\mathcal{O}}^{in}=\frac{\delta\mathcal{O}^{\dagger in}+\delta\mathcal{O}^{in}}{\sqrt{2}}$, $\delta Y_{\mathcal{O}}^{in}=i\frac{\delta\mathcal{O}^{\dagger in}-\delta\mathcal{O}^{in}}{\sqrt{2}}$. This enables us to derive the set of equations describing the quadrature dynamics of our system as,
$\dot{u}={\rm M}u+z^{in}.$ (12)
Here ${u}=(\delta X_{a_{1}},\delta Y_{a_{1}},\delta X_{b_{a}},\delta Y_{b_{a}},\delta X_{b_{m}},\delta Y_{b_{m}})^{T}$, $z^{in}=(\sqrt{\kappa}X_{a_{1}}^{in},\sqrt{\kappa}Y_{a_{1}}^{in},\sqrt{\gamma_{a}}X_{b_{a}}^{in},\sqrt{\gamma_{a}}Y_{b_{a}}^{in},\sqrt{\gamma_{m}}X_{b_{m}}^{in},\sqrt{\gamma_{m}}Y_{b_{m}}^{in},)^{T}$ and the matrix $\rm M$ is given by,
${\rm M}=\begin{pmatrix}-\frac{\kappa_{1}}{2}&-\tilde{\Delta}&0&-G_{a}&0&0\\ \tilde{\Delta}&-\frac{\kappa_{1}}{2}&G_{a}&0&2G_{m}^{{}^{\prime}}&0\\ 0&-G_{a}&-\frac{\gamma_{a}}{2}&\Delta_{a}&J^{\prime}_{m}\sin{\theta}&J^{\prime}_{m}\cos{\theta}\\ G_{a}&0&-\Delta_{a}&-\frac{\gamma_{a}}{2}&-J^{\prime}_{m}\cos{\theta}&J^{\prime}_{m}\sin{\theta}\\ 0&0&-J^{\prime}_{m}\sin{\theta}&J^{\prime}_{m}\cos{\theta}&-\frac{\gamma_{m}}{2}&\omega_{m}^{{}^{\prime}}\\ 2G_{m}^{{}^{\prime}}&0&-J^{\prime}_{m}\cos{\theta}&-J^{\prime}_{m}\sin{\theta}&-\omega_{m}^{{}^{\prime}}&-\frac{\gamma_{m}}{2}.\end{pmatrix},$ (13)
Here the effective couplings $G_{m}^{{}^{\prime}}$ and $G_{a}$ have been assumed to be real for simplicity.